# ReGNet: Reciprocal Space-Aware Long-Range Modeling for Crystalline Property Prediction

### v2

@@ C.2 Related Works on Multi-property prediction
ReGNet-MT 35.0 0.122 36.0 0.21 69.2
ReGNet-LS (3 tasks) 37.0 0.127 37.3 N/A N/A
ReGNet-LS (5 tasks) 49.9 0.136 53.0 0.24 92.1
@@ C.3 MTL and MoE
Multi-Task Learning (MTL) typically employs a shared backbone network to extract common features, followed by task-specific output heads (decoders) that specialize in individual predictions Xiao et al. (2024). In this context, the Mixture of Experts (MoE) framework offers a flexible and scalable approach to capture both shared and task-specific features by leveraging task-relevant experts. MoE has gained significant attention in multi-task scenarios across computer vision and natural language processing due to its adaptability and efficiency Riquelme et al. (2021). An MoE layer consists of a set of expert networks, denoted as $E_{i},\forall i\in[1,N]$, where $N$ is the number of experts, along with a routing network $R$. The output of an MoE layer is the weighted sum of the output $E_{n}(x)$ from every expert, where weights $G_{n}(x)$ are calculated by the routing network $R$ and $x$ represents the model input. Formally, the output of a MoE layer is given by
$y=\sum_{i=1}^{N}G_{n}(x)E_{n}(x).$
The routing network utilizes a noisy Top-$K$ selection mechanism Shazeer et al. (2017) to model the probabilities of each expert and select the top $K$ candidates. This process is defined as:
$G(x)=\text{TopK}(\text{Softmax}(xW+\mathcal{N}(0,1)\text{Softplus}(xW_{\text{noise}}))),$
@@ C.4 Multi-Property Prediction Decoder
ReGNet-MT 24.50 0.218 0.0391 0.0668
ReGNet-LS 39.10 0.221 0.0404 0.0713
For multi-property prediction, we employ ReGNet-MT, a multi-task learning variant of ReGNet in the decoder that incorporates mixture-of-experts (MoE) layers followed by task-specific fully connected heads. We assume that the aggregated node features from the final ReGNet block encapsulate both short-range and long-range information of materials, which can be shared among all property prediction tasks. However, different properties require customized features before individual task heads. Therefore, experts are expected to learn distinct features, while the routing network can determine the optimal expert combination. Unlike the post-processing of the predicted property values after decoders with physical laws Ren et al. (2024), this approach offers more flexibility. Meanwhile, similar properties could have a large expert selection overlap, such that the knowledge can be shared among them. Notably, this expert-sharing mechanism could facilitate positive transfer.
@@ C.5 Detailed Results on Multi-Property Prediction
Figure 5: MoE experts frequencies for each property on the JARVIS dataset and property similarities. The expert selections in the OPT bandgap and MBJ bandgap show the highest similarity, and the positive transfer happens between these two predictions.
In this section, we provide detailed results of using ReGNet-MT on JARVIS and The Material Project datasets. We compare our method with the most accepted MTL method, linear scalarization (LS), and the results are shown in Figure 5, Table 12, and Table 13.
Performance degradation in MTL compared to STL is a well-known phenomenon in the MTL literature. This issue is commonly attributed to the gradient conflict, which arises when gradients from different tasks (properties) point in conflicting directions or have significantly different magnitudes Xiao et al. (2024). Therefore, it is realistic to expect a performance drop in both Table 12 and Table 13 for most properties, and this phenomenon becomes severe when the number of properties increases. However, our ReGNet-MT still matches Matformer and ALIGNN on both datasets.
Surprisingly, ReGNet-MT achieves the best performance for two bandgaps, though some degradation happens. Notably, bandgap (MBJ) surpasses the state-of-the-art (SOTA) results of single-task models, while bandgap (OPT) matches the SOTA performance. To evaluate this improvement, we track the selection frequencies of the 15 experts and visualize the expert selection probabilities for each task (see Figure 5, left). We also compute the cosine similarity among properties via expert selection probabilities (see Figure 5, right). Among the five tasks, two bandgap properties exhibit the highest similarity score of 0.85, indicating a big overlap in the expert selection, which facilitates the positive transfer. This finding explains why ReGNet-MT outperforms single-task training for two bandgaps.

### v3

@@ B.6 Additional Experimental Comparisons with Reciprocal Space-Based Methods
ReciNet-MT 35.0 0.122 36.0 0.21 69.2
ReciNet-LS (3 tasks) 37.0 0.127 37.3 N/A N/A
ReciNet-LS (5 tasks) 49.9 0.136 53.0 0.24 92.1
B.7 Training Time Scale Efficiency Analysis
[…]
Appendix C Multi-task Learning
C.1 Motivation
While our novel usage of reciprocal space representations effectively captures long-range interactions and global periodicity, it still adheres to the traditional approach of predicting properties individually (Schütt et al., 2017; Choudhary & DeCost, 2021; Yan et al., 2024). This approach becomes computationally and memory inefficient when repeatedly applied to multi-property predictions. To address this limitation, Multi-Task Learning (MTL) offers a promising solution by enabling the prediction of multiple properties within a single model (Evgeniou & Pontil, 2004). Additionally, MTL can leverage positive transfer, where knowledge and representations learned from one task can benefit other tasks. This advantage is particularly relevant for materials, as similar properties (e.g., OPT bandgap and MBJ bandgap) are often highly correlated and share similar structural features. However, to our knowledge, few studies have specifically investigated multi-property prediction for materials.
C.2 Related Works on Multi-property prediction
@@ C.3 MTL and MoE
Multi-Task Learning (MTL) typically employs a shared backbone network to extract common features, followed by task-specific output heads (decoders) that specialize in individual predictions (Xiao et al., 2024). In this context, the Mixture of Experts (MoE) framework offers a flexible and scalable approach to capture both shared and task-specific features by leveraging task-relevant experts. MoE has gained significant attention in multi-task scenarios across computer vision and natural language processing due to its adaptability and efficiency (Riquelme et al., 2021). An MoE layer consists of a set of expert networks, denoted as $E_{i},\forall i\in[1,N]$, where $N$ is the number of experts, along with a routing network $R$. The output of an MoE layer is the weighted sum of the output $E_{n}(x)$ from every expert, where weights $G_{n}(x)$ are calculated by the routing network $R$ and $x$ represents the model input. Formally, the output of a MoE layer is given by
$y=\sum_{i=1}^{N}G_{n}(x)E_{n}(x).$ (26)
The routing network utilizes a noisy Top-$K$ selection mechanism (Shazeer et al., 2017) to model the probabilities of each expert and select the top $K$ candidates. This process is defined as:
$G(x)=\text{TopK}(\text{Softmax}(xW+\mathcal{N}(0,1)\text{Softplus}(xW_{\text{noise}}))),$ (27)
@@ C.4 Multi-Property Prediction Decoder
ReciNet-MT 24.50 0.218 0.0391 0.0668
ReciNet-LS 39.10 0.221 0.0404 0.0713
For multi-property prediction, we employ ReciNet-MT, a multi-task learning variant of ReciNet in the decoder that incorporates mixture-of-experts (MoE) layers followed by task-specific fully connected heads. We assume that the aggregated node features from the final ReciNet block encapsulate both short-range and long-range information of materials, which can be shared among all property prediction tasks. However, different properties require customized features for individual task heads. Therefore, experts are expected to learn distinct features, while the routing network can determine the optimal expert combination. Unlike the post-processing of the predicted property values after decoders with physical laws (Ren et al., 2024), this approach offers more flexibility. Meanwhile, similar properties could have a large expert selection overlap, such that the knowledge can be shared among them. Notably, this expert-sharing mechanism could facilitate positive transfer.
@@ C.5 Detailed Results on Multi-Property Prediction
Figure 4: MoE experts’ frequencies for each property on the JARVIS dataset and property similarities. The expert selections in the OPT bandgap and MBJ bandgap show the highest similarity, and the positive transfer happens between these two predictions.
In this section, we provide detailed results of using ReciNet-MT on JARVIS and The Materials Project datasets. We compare our method with the most accepted MTL method, linear scalarization (LS), and the results are shown in Figure˜4, Table˜13, and Table˜14.
Performance degradation in MTL compared to STL is a well-known phenomenon in the MTL literature. This issue is commonly attributed to the gradient conflict, which arises when gradients from different tasks (properties) point in conflicting directions or have significantly different magnitudes (Xiao et al., 2024). Therefore, it is realistic to expect a performance drop in both Table˜13 and Table˜14 for most properties, and this phenomenon becomes severe when the number of properties increases. However, our ReciNet-MT still matches Matformer and ALIGNN on both datasets.
Surprisingly, ReciNet-MT achieves the best performance for two bandgaps, though some degradation happens. Notably, bandgap (MBJ) surpasses the state-of-the-art (SOTA) results of single-task models, while bandgap (OPT) matches the SOTA performance. To evaluate this improvement, we track the selection frequencies of the 15 experts and visualize the expert selection probabilities for each task (see Figure˜4, left). We also compute the cosine similarity among properties via expert selection probabilities (see Figure˜4, right). Among the five tasks, two bandgap properties exhibit the highest similarity score of 0.85, indicating a big overlap in the expert selection, which facilitates the positive transfer. This performance gain can be attributed to positive transfer across related properties; specifically, both band gap (OPT) and band gap (MBJ) are governed by the underlying electronic structure, enabling shared representations to benefit both tasks. This finding explains why ReciNet-MT outperforms single-task training for two band gaps.